Assume you are the CFO of AIFS. Your analyst reports the following information (Use the following information for the remainder of the assignment):

  • Current exchange rate is $1.18/€.
  • Forward rate is $1.195/€.
  • Expected final sales volume is 30,000. Worst case scenario is volume of 20,000. Best case scenario is volume of 40,000.
  • Cost per student is €2500.
  • Option premium is 1.2% of USD strike price.
  • Option strike price is $1.17/€.

4. Using the above information

a) What is the total projected costs (for all three scenarios) in dollars at the current exchange rate?

b) What are the total costs (for all three scenarios) if you use a forward contract to hedge?

c) What is the total option premium for each scenario?

5. As the CFO, you decided not to hedge. Assuming expected final sales volume is 30,000, what are your total costs

a) if the exchange rate remains at $1.18/€? Let’s call this the baseline scenario.

b) if the exchange rate will be $1.3/€? How does this compare to the baseline case?

c) if the exchange rate will be $1.1/€? How does this compare to the baseline case?

6. As the CFO, you decided to hedge using forward contracts. Assume that the expected final sales volume is 30,000. What are your total benefit/cost and the percentage benefit/cost from hedging (compared to no hedging)

a) if the exchange rate remains at $1.18/€?

b) if the exchange rate will be $1.3/€?

c) if the exchange rate will be $1.1/€?

7. As the CFO, you decided to hedge using option contracts. Assuming expected final sales volume is 30,000, what are your total benefit/cost and the percentage benefit/cost from hedging (compared to no hedging)

a) if the exchange rate remains at $1.18/€?

b) if the exchange rate will be $1.3/€?

c) if the exchange rate will be $1.1/€?

8. What is the most profitable strategy for the case in which the expected final sales volume is 30,000 (no hedge, forward contract, or option contract)

a) if the exchange rate remains at $1.18/€?

b) if the exchange rate will be $1.3/€?

c) if the exchange rate will be $1.1/€?

d) Is there a best strategy? Why? Note that you only need to provide a few sentences about which strategy to use from no hedge, forward contract, or option contract strategies.
Answer #1
4a. Expected sales worst case best case
units 30000 20000 40000
cost per student ( in euro) €                 2,500 €                 2,500 €              2,500
current exchange rate ($/euro) 1.18 1.18 1.18
so the cost in terms of $ = cost in euro* exhange rate 2500*1.18= $                 2,950 $                 2,950 $              2,950
total cost= cost per student * units =30000*2950= $       88,500,000 $       59,000,000 $ 118,000,000
4b. Expected sales worst case best case
units 30000 20000 40000
cost per student ( in euro) €                 2,500 €                 2,500 €              2,500
forward cobtract exchange rate ($/euro) 1.195 1.195 1.195
so the cost in terms of $ = cost in euro* exhange rate 2500*1.195= $                 2,988 $                 2,988 $              2,988
total cost= cost per student * units =30000*2998= $       89,625,000 $       59,750,000 $ 119,500,000
4c. Expected sales worst case best case
units 30000 20000 40000
cost per student ( in euro) €                 2,500 €                 2,500 €              2,500
option strike price ($/euro) 1.17 1.17 1.17
Option Premium 1.20% 1.20% 1.20%
Premium 1.17*1.2%                      0.014                      0.014                   0.014
Premium paid in each senario($/ euro) 30000*0.014 ans so on                   421.20                   280.80                561.60
total Premium cost= premium per student * units =2500*421.2= $         1,053,000 $             702,000 $      1,404,000
5 exchange rate ($/euro) 1.18 1.3 1.1
sales In units 30000 30000 30000
cost per student ( in euro) €                2,500 €                 2,500 €                 2,500
cost per student (in $)= exchange rate * cost per student (in euro) =2500*1.18= $2950 2500*1.3= $3250 2500*1.1= $2750
total cot of student = sold units* cost per student (in $) 30000*2950= $88,500,000 30000*3250= $ 97,500,000 30000*2750= $82,500,000
Compare Equal to baseline Profitable than baseline loss as compared to baseline